OpenStudy (anonymous):

1. Write a related equation. w ÷ 9 = 14 (Points : 1) w = 14 ÷ 9 w = 14 9 w = 14 9 w = 14 + 9

OpenStudy (anonymous):

how do you undo division?

OpenStudy (anonymous):

what is its opposite operation?

OpenStudy (anonymous):

not a clue

OpenStudy (anonymous):

what is addition's opposite operation?

OpenStudy (igreen):

You have to do the opposite to both sides in order to keep the balance of an equation/inequality.

OpenStudy (igreen):

So what's the opposite of dividing? @cheyycheyy1234

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

there you go... in fact, multiplying by a number, say n and multiplying by its reciprocal \(\large{\frac{1}{n}}\) are opposite operations. we call the latter one division.

OpenStudy (igreen):

Okay, so you have to multiply 9 to both sides..so which option is it? @cheyycheyy1234

OpenStudy (anonymous):

im confused

OpenStudy (igreen):

You want to get 'w' by itself. To keep the balance of the equation, you have to do the opposite. 9 is being divide by 'w'. To get 'w' by itself you have to do the opposite to the other side of the equation, which is multiplying 9.

OpenStudy (igreen):

So which option has 9 being multiplied?

OpenStudy (igreen):

Option B and C look the same to me.

OpenStudy (anonymous):

ok now i gotcha thanks guys

OpenStudy (igreen):

If any of the options look like w = 14 * 9, w = 14 x 9, w = 14(9), or w = 14 · 9 they're correct.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @cheyycheyy1234